Grade 12th passMechanics

A particle moves in a straight line and its position x at the time t is given by x^=2+t its acceleration is given by

Prakash
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
X^2 = 2 + t
d(x^2) / dv = d(2 + t)/dv
2dx/dv = dt/dv
2xdx = dt
dx/dt = 1/(2x)
v = 1/(2x)

Differentiating above equation
dv/dx = -1/(2x^2)

Acceleration is given by
a = vdv/dx
= 1/(2x) * -1/(2x^2)
= -1/(4x^3)

It’s acceleration is given by -1/(4x^3)